# Complementary Probes of Warped Extra Dimension: Colliders, Gravitational Waves and Primordial Black Holes from Phase Transitions

### v1

@@ Abstract
Extra Dimensions, Phase Transitions, Gravitational Waves, Dark Matter, Black Holes
pacs:
12.60.-i, 04.70.-s, 04.20.-q, 98.80.-k, 64.60.-i
I Introduction
@@
The detection of a primordial SGWB would provide a unique probe of the early Universe. Primordial GWs travel from the early Universe to the present day with negligible (Planck scale suppressed) interaction, unlike other cosmic relics such as photons and neutrinos. This long free path ensures that the source information carried by primoridial GWs remains uncontaminated. So far, the LVK collaborations have searched for SGWBs and found none, setting an upper limit on this class of signals at kHz frequencies Abbott et al. (2017, 2019, 2021). At nHz frequencies, instead, the Pulsar Timing Arrays (PTAs) collaborations Afzal et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023) have reported strong evidence of a SGWB, which they can ascribe to astrophysical and/or cosmological sources Agazie et al. (2023); Afzal et al. (2023); Antoniadis et al. (2024). Forthcoming and future experiments Weltman et al. (2020); Garcia-Bellido et al. (2021); Abe et al. (2021); Badurina et al. (2020); El-Neaj et al. (2020); Sesana et al. (2021); Amaro-Seoane et al. (2017); Luo et al. (2016); Ruan et al. (2020); Kawamura et al. (2021); Corbin and Cornish (2006); Punturo et al. (2010); Reitze et al. (2019); Aggarwal et al. (2021); Berlin et al. (2022); Herman et al. (2023); Bringmann et al. (2023); Valero et al. (2024) will improve current SGWB measurements by increasing sensitivities around nHz and kHz frequencies as well as extending them to the entire nHz-GHz frequency range.
@@
It is well known that spontaneous symmetry breakings occurring in the early universe may undergo strong first-order phase transitions (FOPTs). During FOPTs, bubbles containing the broken phase nucleate, expand, collide, and percolate. The bubble collisions, along with the motion of the surrounding thermal plasma, are violent, energetic processes that generate GWs Witten (1984); Hogan (1986). The overall signal is a SGWB with a distinct frequency shape, peaking at the frequency $f_{\rm p}\sim T_{0}T/M_{\rm Pl}$ (see Ref. Auclair et al. (2023) for a review). In particular, the SGWB from a FOPT at the electroweak (EW) scale peaks at mHz frequencies, within the LISA band, while the SGWB from a FOPT occurring at the EeV scale peaks at kHz frequencies, i.e., in the frequency band the ground-based GW interferometers are sensitive to. Notably, none of the spontaneous symmetry breakings predicted by the Standard Model (SM) occur via a FOPT Kajantie et al. (1996); Laine and Meyer (2015); Bhattacharya et al. (2014), making the potential detection of a FOPT SGWB a revolutionary target. Its detection would indeed be striking proof of the existence of beyond the SM (BSM) physics.
@@
BSM physics is expected for several reasons. The SM suffers from severe drawbacks, including the dark matter (DM) puzzle, the baryon asymmetry of the universe, and the electroweak hierarchy problem. The Randall-Sundrum (RS) setup Randall and Sundrum (1999) is a promising option to solve these problems. This theory involves a mechanism that breaks the conformal symmetry by generating a potential for the radion, the degree of freedom associated with the distance between the electroweak (EW) and Planck branes along the warped space dimension. If this mechanism is realized à la Goldberger-Wise Goldberger and Wise (1999), the conformal symmetry breaking due to the radion happens via a FOPT Creminelli et al. (2002); Randall and Servant (2007); Kaplan et al. (2006); Nardini et al. (2007); Hassanain et al. (2007); Konstandin et al. (2010); Bunk et al. (2018); Dillon et al. (2018); von Harling and Servant (2018); Megías et al. (2018), and the induced warped factor naturally explains the EW-Planck hierarchy. Such a radion FOPT is a key ingredient for solving the baryon asymmetry in RS-inspired models Nardini et al. (2007); Konstandin and Servant (2011); Bruggisser et al. (2023, 2018). Moreover, it may help address the DM puzzle in RS-like models, as we investigate in the present work for the first time. The idea is that the radion FOPT can trigger the formation of primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974) with masses and abundances such that PBHs mimic DM while being compatible with current DM and PBH constraints. For instance, the PBH-to-DM density ratio $f_{{\rm PBH}}\equiv\rho_{{\rm PBH}}/\rho_{{\rm DM}}\simeq 1$ is experimentally allowed when the PBHs have a monochromatic mass distribution in the range Carr and Kuhnel (2022); Bagui et al. (2023)
@@
Various mechanisms acting during first-order phase transitions lead to the density inhomogeneities required for PBH formation Hawking et al. (1982); Crawford and Schramm (1982); Kodama et al. (1982); Hsu (1990); Moss (1994); Khlopov et al. (1998); Lewicki and Vaskonen (2020); Liu et al. (2022); Hashino et al. (2022); Gross et al. (2021); Baker et al. (2021); Kawana and Xie (2022); He et al. (2024); Hashino et al. (2023); Kawana et al. (2023); Lewicki et al. (2023); Gouttenoire and Volansky (2024); Salvio (2023). One such mechanism unavoidably arises in slow, extremely supercooled FOPTs Sato et al. (1981); Kodama et al. (1981); Maeda et al. (1982); Sato et al. (1982); Hawking et al. (1982); Kodama et al. (1982); Lewicki and Vaskonen (2020); Ashoorioon et al. (2021); Kawana and Xie (2022); Liu et al. (2022); Jung and Okui (2021); Hashino et al. (2023); Huang and Xie (2022); Kawana et al. (2022, 2023); Kierkla et al. (2023, 2024); Hashino et al. (2022); He et al. (2024); Gehrman et al. (2023); Lewicki et al. (2023); Gouttenoire and Volansky (2024); Gouttenoire (2024); Baldes and Olea-Romacho (2024); Salvio (2023, 2024); Banerjee and Dey (2024); Gonçalves et al. (2024); Conaci et al. (2024); Cai et al. (2024); Arteaga et al. (2024). Its rationale is rather intuitive. When the FOPT is extremely supercooled, the temperature of the thermal bath in the unbroken phase can be so low that the radiation energy density is negligible compared to the vacuum energy density. The FOPT from this vacuum-dominated phase to the broken one thus occurs with huge reheating. Indeed, in the approximation that reheating is instantaneous, the volume occupied by a just-nucleated bubble (full of broken phase) contains the same energy density immediately before and after the bubble nucleation, which is possible only if a large radiation energy density appears after nucleation. In other words, at the time of nucleation, the broken-phase energy density inside the bubble is equal to that of the unbroken phase where the bubble has appeared. However, this energy density equality no longer holds at later times: inside the bubble, the energy density dilutes over time as radiation, while it remains constant in the unbroken phase where the vacuum energy dominates. This implies that, at a given time during the FOPT, an older bubble has less energy density than a younger one, and the larger their nucleation-time gap, the greater their energy density difference. For this reason, the energy density inhomogeneities are more pronounced if the FOPT is slow, i.e., the times at which the FOPT starts and ends are well separated.
@@
At a quantitative level, the aforementioned PBH formation mechanism is not settled. For the PBH mass distribution, we consider the leading-order contribution yielding a monochromatic mass spectrum. Nevertheless, subleading corrections to this result are plausible Lewicki et al. (2023). Moreover, general relativistic numerical simulations of PBH formation Shibata and Sasaki (1999); Musco et al. (2005); Harada et al. (2013); Musco (2019); Musco et al. (2021); Germani and Musco (2019); Musco (2019); Escrivà et al. (2020); Escrivà and Romano (2021); Escrivà et al. (2023) do not include the bubble reheating processes, but our estimate of $\mathcal{O}(1)$ overdensities leads us to assume that, to some extent, the collapse should occur. Given these uncertainties, we follow the available established estimates Liu et al. (2022); Hashino et al. (2022); He et al. (2024); Hashino et al. (2023); Kawana et al. (2023); Gouttenoire and Volansky (2024); Salvio (2023), with the understanding that they will need to be replaced with more precise ones when available. In any case, future improvements should not revolutionize the qualitative conclusions we achieve in the present study.
@@
In models of warped extra dimensions, the radion naturally undergoes a supercooled FOPT Baratella et al. (2019); Megías et al. (2020, 2023). In this work, we analyze the phenomenological implications of this FOPT, including the aforementioned PBH formation. For concreteness, we focus on two simple warped setups involving either two or three branes, with the SM fields appropriately localized on one of them. We consider either one or the other scenario depending on the radion breaking scale $\rho$. Specifically, for scenarios where $\rho$ is at TeV or above, we assume the setup with the SM fields localized on the IR brane, as in the original RS scenario Randall and Sundrum (1999) [^1] In more complicated (realistic) theories, the SM fields, except possibly the Higgs boson and the right-handed top quark, propagate in the bulk, and a realistic theory of flavor can be constructed. For simplicity, and to stress the main points of our mechanism for the production of PBH and GWs, we assume here that all the SM fields are four-dimensional and propagate on a brane. The flavor problem then remains as in the SM. Moreover, the precision electroweak data are not in tension with to Kaluza-Klein (KK) states, as we shall see later on.. For scenarios with $\rho\lesssim 1$ TeV, we instead work with the the SM fields localised on an additional, intermediate brane (between the IR and UV branes) with scale $\rho_{T}\sim 1$ TeV. In both setups, the only degrees of freedom propagating in the bulk are the graviton and the radion, with well-defined couplings to the SM brane fields. Additionally, the warped factors between the branes naturally explain the branes’ scale hierarchies. In both setups, the Goldberger-Wise stabilization mechanism is assumed.
@@
Once the extra dimension is integrated out, our warped setups contain towers of heavy states, the KK modes of all particles propagating in the bulk. They also contain the radion, which typically remains the lightest BSM state and plays a relevant role in collider and early-universe phenomenology. On the one hand, the radion, together with the KK particles, could give rise to resonant signals at the LHC or future colliders, whose so-far null searches impose upper bounds on the couplings between the SM particles and the radion and KK fields Arganda et al. (2024). On the other hand, as previously explained, the radion undergoes a FOPT that potentially generates a detectable SGWB and a PBH population solving the DM puzzle. The overall phenomenology is then very rich and, as we will conclude in the present studies, can be tested by fully leveraging the synergy of present and future GW, PBH, and collider experiments.
@@
The paper is organized as follows. In Sec. II we introduce the two warped extra-dimension setups that we investigate. In Sec. III we study the radion FOPT and the key temperatures that characterize it. Section IV provides the relevant $\ae$ for PBH formation and recasts the main PBH observables in terms of the parameters of our warped models. Section V quantifies the SGWB produced by the radion FOPT and its detection prospects at present and future GW experiments. In Sec. VI we investigate the models’ collider signatures due to the couplings between the SM particles, the graviton, and the radion, and we study some astrophysical bounds that constrain the radion lifetime. In Sec. VII we compare the parameter reach of the PBH, GW, and collider searches obtained in the previous sections, highlight their synergy and complementarity, and comment on our main results. Finally, in App. A we derive the $\ae$ for the relevant FOPT temperatures obtained in the thick-wall approximation, while in App. B we show how to improve them through a semi-analytical approach.
II Two-model setups
@@
II.1 High-energy setup: $\rho\gtrsim$ TeV
@@
To stabilize the brane distance, $\log(z_{1}/z_{0})$, we must break the conformal symmetry. We do it by introducing a heavy scalar field $\phi$ that propagates in the bulk, i.e. the spacetime between the branes. This field has the bulk potential $V(\phi)$ and brane potentials
@@
We obtain both $V(\phi)$ and $A(z)$ by solving the 5D Einstein equations with the so-called superpotential method DeWolfe et al. (2000). We consider the superpotential $W(\phi)=6kM_{5}^{3}+uk\phi^{2}$, where $u>0$ is a dimensionless parameter controlling the backreaction of the scalar field on the 5D metric, while $M_{5}$ is the 5D Planck scale. We also fix the brane tensions $\Lambda_{a}$ in Eq. (3) to $\Lambda_{0}=-W(v_{0})$ and $\Lambda_{1}=-W(v_{1})+12kM_{5}^{3}\lambda_{1}$, with $\lambda_{1}<0$ being a free detuning parameter Goldberger and Wise (1999). Such a superpotential and boundary conditions are known to yield the right AdS5 UV behavior in the $u\ll 1$ regime DeWolfe et al. (2000).
By breaking the conformal symmetry, the heavy field $\phi$ induces an effective potential $V_{{\rm eff}}$ to the radion field $\chi(x)$, the degree of freedom describing the excitation $\tilde{\chi}(x)=\chi(x)-\rho$ around the location of the $\mathcal{B}_{1}$ brane. At the minimum of the potential, the radion acquires the vacuum expectation value $\langle\chi\rangle=\rho$, thus fixing the brane distance by $z_{1}/z_{0}=k/\rho$, fixed by the condition $v_{1}=v_{0}(z_{1}/z_{0})^{u}$. Moreover, the second derivative of the potential at the minimum permits to estimate the radion mass $m_{\chi}$. In the small-backreaction regime, such a mass can be approximated as Megías et al. (2023)
@@
On the other hand, at finite temperature, a second metric background is also compatible with the 5D Einstein equations when $A(z)\sim\log z$ at small $z$ Creminelli et al. (2002). This solution is the 5D Anti-de-Sitter Schwarzschild (AdS-S) metric which describes the spacetime of a black hole (BH) with event horizon at $z=z_{h}$. In this second metric background, the conformal invariance is unbroken and, in the duality language, this solution is characterized by SU(N) strongly-coupled (massless) gauge fields that are deconfined. The AdS/CFT correspondence ralates $N$ to $k$ through the equality
@@
where $M_{P}=2.4\times 10^{15}$ TeV is the 4D reduced Planck scale. Requiring the 5D gravity theory to be perturbative imposes the bound $N\gtrsim 5$ Agashe et al. (2007); Chen et al. (2014).
The fact that both the AdS-S and warped metric solutions coexist implies the existence of two spacetime background phases. In the regime of small backreaction $u\ll 1$ and tiny temperature, the energy gap between the two phases is dominated by the difference $E_{0}=\left|V_{{\rm eff}}(0)-V_{{\rm eff}}(\rho)\right|$ which can be approximated as Megías et al. (2020)
@@
Moreover, the free energy associated to these two phases presents a barrier between them, implying the existence of a phase transition between the AdS and RS phases Creminelli et al. (2002). In the dual language, such a phase transition corresponds to a deconfined-confined FOPT.
@@
II.2 Low-energy setup: $\rho\lesssim$ TeV

### v2

@@ Abstract
1 Introduction
@@
The detection of a primordial SGWB would provide a unique probe of the early Universe. Primordial GWs travel from the early Universe to the present day with negligible (Planck scale suppressed) interaction, unlike other cosmic relics such as photons and neutrinos. This long free path ensures that the source information carried by primordial GWs remains uncontaminated. So far, the LVK collaborations have searched for SGWBs and found none, setting an upper limit on this class of signals at kHz frequencies LIGOScientific:2016jlg ; LIGOScientific:2019vic ; KAGRA:2021kbb . At nHz frequencies, instead, the Pulsar Timing Arrays (PTAs) collaborations NANOGrav:2023hvm ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog have reported strong evidence of a SGWB, which they can ascribe to astrophysical and/or cosmological sources NANOGrav:2023hfp ; NANOGrav:2023hvm ; EPTA:2023xxk . Forthcoming and future experiments Weltman:2018zrl ; Garcia-Bellido:2021zgu ; MAGIS-100:2021etm ; Badurina:2019hst ; AEDGE:2019nxb ; Sesana:2019vho ; LISA:2017pwj ; TianQin:2015yph ; Ruan:2018tsw ; Kawamura:2020pcg ; Corbin:2005ny ; Punturo:2010zz ; Reitze:2019iox ; Aggarwal:2020olq ; Berlin:2021txa ; Herman:2022fau ; Bringmann:2023gba ; Valero:2024ncz will improve current SGWB measurements by increasing sensitivities around nHz and kHz frequencies as well as extending them to the entire nHz-GHz frequency range.
@@
It is well known that spontaneous symmetry breakings occurring in the early universe may undergo strong first-order phase transitions (FOPTs). During FOPTs, bubbles containing the broken phase nucleate, expand, collide, and percolate. The bubble collisions, along with the motion of the surrounding thermal plasma, are violent, energetic processes that generate GWs Witten:1984rs ; Hogan:1986qda . The overall signal is a SGWB with a distinct frequency shape, peaking at the frequency $f_{\rm p}\sim T_{0}T/M_{\rm Pl}$ (see Ref. LISACosmologyWorkingGroup:2022jok for a review). In particular, the SGWB from a FOPT at the electroweak (EW) scale peaks at mHz frequencies, within the LISA band, while the SGWB from a FOPT occurring at the EeV scale peaks at kHz frequencies, i.e., in the frequency band the ground-based GW interferometers are sensitive to. Notably, none of the spontaneous symmetry breakings predicted by the Standard Model (SM) occur via a FOPT Kajantie:1996mn ; Laine:2015kra ; Bhattacharya:2014ara , making the potential detection of a FOPT SGWB a revolutionary target. Its detection would indeed be striking proof of the existence of beyond the SM (BSM) physics.
@@
BSM physics is expected for several reasons. The SM suffers from severe drawbacks, including the dark matter (DM) puzzle, the baryon asymmetry of the universe, and the electroweak hierarchy problem. The Randall-Sundrum (RS) setup Randall:1999ee is a promising option to solve these problems. This theory involves a mechanism that breaks the conformal symmetry by generating a potential for the radion, the degree of freedom associated with the distance between the electroweak (EW) and Planck branes along the warped space dimension. If this mechanism is realized à la Goldberger-Wise Goldberger:1999uk , the conformal symmetry breaking due to the radion happens via a FOPT Creminelli:2001th ; Randall:2006py ; Kaplan:2006yi ; Nardini:2007me ; Hassanain:2007js ; Konstandin:2010cd ; Bunk:2017fic ; Dillon:2017ctw ; vonHarling:2017yew ; Baratella:2018pxi ; Megias:2018sxv ; Agashe:2019lhy ; Fujikura:2019oyi ; Megias:2020vek ; Bigazzi:2020phm ; Agashe:2020lfz ; Megias:2023kiy ; Mishra:2023kiu ; Mishra:2024ehr ; Barbosa:2024pyn ; Agrawal:2025wvf ; Gherghetta:2025krk , and the induced warped factor naturally explains the EW-Planck hierarchy. Such a radion FOPT is a key ingredient for solving the baryon asymmetry in RS-inspired models Nardini:2007me ; Konstandin:2011ds ; Bruggisser:2022rdm ; Bruggisser:2018mus . Moreover, it may help address the DM puzzle in RS-like models, as we investigate in the present work for the first time. The idea is that the radion FOPT can trigger the formation of primordial black holes (PBHs) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx with masses and abundances such that PBHs mimic DM while being compatible with current DM and PBH constraints. For instance, the PBH-to-DM density ratio $f_{{\rm PBH}}\equiv\rho_{{\rm PBH}}/\rho_{{\rm DM}}\simeq 1$ is experimentally allowed when the PBHs have a monochromatic mass distribution in the range Carr:2021bzv ; LISACosmologyWorkingGroup:2023njw
@@
Various mechanisms acting during FOPTs lead to the density inhomogeneities required for PBH formation Hawking:1982ga ; Crawford:1982yz ; Kodama:1982sf ; Hsu:1990fg ; Moss:1994iq ; Khlopov:1998nm ; Lewicki:2019gmv ; Liu:2021svg ; Hashino:2021qoq ; Gross:2021qgx ; Baker:2021sno ; Kawana:2021tde ; He:2022amv ; Hashino:2022tcs ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Salvio:2023ynn ; Ai:2024cka . One such mechanism unavoidably arises in slow, extremely supercooled FOPTs Sato:1981bf ; Kodama:1981gu ; Maeda:1981gw ; Sato:1981gv ; Hawking:1982ga ; Kodama:1982sf ; Lewicki:2019gmv ; Ashoorioon:2020hln ; Kawana:2021tde ; Liu:2021svg ; Jung:2021mku ; Hashino:2022tcs ; Huang:2022him ; Kawana:2022lba ; Kawana:2022olo ; Kierkla:2022odc ; Kierkla:2023von ; Hashino:2021qoq ; He:2022amv ; Gehrman:2023esa ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2023qya ; Goncalves:2024vkj ; Conaci:2024tlc ; Cai:2024nln ; Arteaga:2024vde . Its rationale is rather intuitive. When the FOPT is extremely supercooled, the temperature of the thermal bath in the unbroken phase can be so low that the radiation energy density is negligible compared to the vacuum energy density. The FOPT from this vacuum-dominated phase to the broken one thus occurs with huge reheating. Indeed, in the approximation that reheating is instantaneous, the volume occupied by a just-nucleated bubble (full of broken phase) contains the same energy density immediately before and after the bubble nucleation, which is possible only if a large radiation energy density appears after nucleation. In other words, at the time of nucleation, the broken-phase energy density inside the bubble is equal to that of the unbroken phase where the bubble has appeared. However, this energy density equality no longer holds at later times: inside the bubble, the energy density dilutes over time as radiation, while it remains constant in the unbroken phase where the vacuum energy dominates. This implies that, at a given time during the FOPT, an older bubble has less energy density than a younger one, and the larger their nucleation-time gap, the greater their energy density difference. For this reason, the energy density inhomogeneities are more pronounced if the FOPT is slow, i.e., the times at which the FOPT starts and ends are well separated.
@@
At a quantitative level, the aforementioned PBH formation mechanism is not settled. For the PBH mass distribution, we consider the leading-order contribution yielding a monochromatic mass spectrum. Nevertheless, subleading corrections to this result are plausible Lewicki:2023ioy . Moreover, general relativistic numerical simulations of PBH formation Shibata:1999zs ; Musco:2004ak ; Harada:2013epa ; Musco:2018rwt ; Musco:2020jjb ; Germani:2018jgr ; Musco:2018rwt ; Escriva:2019phb ; Escriva:2021pmf ; Escriva:2022bwe do not include the bubble reheating processes, but our estimate of $\mathcal{O}(1)$ overdensities leads us to assume that, to some extent, the collapse should occur. Given these uncertainties, we follow the available established estimates Liu:2021svg ; Hashino:2021qoq ; He:2022amv ; Hashino:2022tcs ; Kawana:2022olo ; Gouttenoire:2023naa ; Salvio:2023ynn , with the understanding that they will need to be replaced with more precise ones when available. In any case, future improvements should not revolutionize the qualitative conclusions we achieve in the present study.
@@
In models of warped extra dimensions, the radion naturally undergoes a supercooled FOPT Baratella:2018pxi ; Megias:2020vek ; Megias:2023kiy . In this work, we analyze the phenomenological implications of this FOPT, including the aforementioned PBH formation. For concreteness, we focus on two simple warped setups involving either two or three branes, with the SM fields appropriately localized on one of them. We consider either one or the other scenario depending on the radion breaking scale $\rho$. Specifically, for scenarios where $\rho$ is at TeV or above, we assume the setup with the SM fields localized on the IR brane, as in the original RS scenario Randall:1999ee  [^1] In more complicated (realistic) theories, the SM fields, except possibly the Higgs boson and the right-handed top quark, propagate in the bulk, and a realistic theory of flavor can be constructed. For simplicity, and to stress the main points of our mechanism for the production of PBH and GWs, we assume here that all the SM fields are four-dimensional and propagate on a brane. The flavor problem then remains as in the SM. Moreover, the precision electroweak data are not in tension with to Kaluza-Klein (KK) states, as we shall see later on.. For scenarios with $\rho\lesssim 1$ TeV, we instead work with the the SM fields localised on an additional, intermediate brane (between the IR and UV branes) with scale $\rho_{T}\sim 1$ TeV. In both setups, the only degrees of freedom propagating in the bulk are the graviton and the radion, with well-defined couplings to the SM brane fields. Additionally, the warped factors between the branes naturally explain the branes’ scale hierarchies. In both setups, the Goldberger-Wise stabilization mechanism is assumed.
@@
Once the extra dimension is integrated out, our warped setups contain towers of heavy states, the KK modes of all particles propagating in the bulk. They also contain the radion, which typically remains the lightest BSM state and plays a relevant role in collider and early-universe phenomenology. On the one hand, the radion, together with the KK particles, could give rise to resonant signals at the LHC or future colliders, whose so-far null searches impose upper bounds on the couplings between the SM particles and the radion and KK fields Arganda:2024nyi . On the other hand, as previously explained, the radion undergoes a FOPT that potentially generates a detectable SGWB and a PBH population solving the DM puzzle. The overall phenomenology is then very rich and, as we will conclude in the present studies, can be tested by fully leveraging the synergy of present and future GW, PBH, and collider experiments.
@@
The paper is organized as follows. In Sec. 2 we introduce the two warped extra-dimension setups that we investigate. In Sec. 3 we study the radion FOPT and the key temperatures that characterize it. Section 4 provides the relevant $\ae$ for PBH formation and recasts the main PBH observables in terms of the parameters of our warped models. Section 5 quantifies the SGWB produced by the radion FOPT and its detection prospects at present and future GW experiments. In Sec. 6 we investigate the models’ collider signatures due to the couplings between the SM particles, the graviton, and the radion, and we study some astrophysical bounds that constrain the radion lifetime. In Sec. 7 we compare the parameter reach of the PBH, GW, and collider searches obtained in the previous sections, highlight their synergy and complementarity, and comment on our main results. Finally, in App. A we derive the $\ae$ for the relevant FOPT temperatures obtained in the thick-wall approximation, while in App. B we show how to improve them through a semi-analytical approach.
2 Two-model setups
@@
2.1 High-energy setup: $\rho\gtrsim$ TeV
@@
To stabilize the brane distance, $\log(z_{1}/z_{0})$, we must break the conformal symmetry. We do it by introducing a (stabilizing) scalar field $\phi$ that propagates in the bulk, i.e. the spacetime between the branes. This field has the bulk potential $V(\phi)$ and brane potentials
@@
We obtain both $V(\phi)$ and $A(z)$ by solving the 5D Einstein equations with the so-called superpotential method DeWolfe:1999cp . We consider the superpotential $W(\phi)=6kM_{5}^{3}+uk\phi^{2}$, where $u>0$ is a dimensionless parameter controlling the backreaction of the scalar field on the 5D metric, while $M_{5}$ is the 5D Planck scale. We also fix the brane tensions $\Lambda_{a}$ in Eq. (3) to $\Lambda_{0}=-W(v_{0})$ and $\Lambda_{1}=-W(v_{1})+12kM_{5}^{3}\lambda_{1}$, with $\lambda_{1}<0$ being a free detuning parameter Goldberger:1999uk . Such a superpotential and boundary conditions are known to yield the right AdS5 UV behavior (i.e. $A(z)\sim\log z$ in the region $z\sim z_{0}$) in the case where the constant $u$ fulfills the condition $u\ll 1$, which means small backreaction DeWolfe:1999cp .
By breaking the conformal symmetry, and evaluating the classical action at the background of the stabilizing field $\phi$, an effective potential $V_{{\rm eff}}$ on the radion field $\chi(x)$ is induced, where $\chi(x)\equiv\rho+\tilde{\chi}(x)$ is the full radion field containing the degree of freedom describing the excitation $\tilde{\chi}(x)=\chi(x)-\rho$ around the location of the $\mathcal{B}_{1}$ brane. At the minimum of the potential, the radion acquires the vacuum expectation value $\langle\chi\rangle=\rho$, thus fixing the brane distance by $z_{1}/z_{0}=k/\rho$, fixed by the condition $v_{1}=v_{0}(z_{1}/z_{0})^{u}$. Moreover, the second derivative of the potential at the minimum permits to estimate the radion mass $m_{\chi}$. In the small-backreaction regime, such a mass can be approximated as Megias:2023kiy
@@
On the other hand, at finite temperature, a second metric background is also compatible with the 5D Einstein equations when $A(z)\sim\log z$ at small $z$ Creminelli:2001th . This solution is the 5D Anti-de-Sitter Schwarzschild (AdS-S) metric which describes the spacetime of a black hole (BH) with event horizon at $z=z_{h}$. As was the case for the AdS5 solution, this metric background is symmetric under conformal invariance, although the latter is explicitly broken by the stabilizing field $\phi$, mainly via the brane potentials which fix the value of $\phi$ at the location of the different branes. This second solution is characterized by SU(N) strongly-coupled (massless) gauge fields that are deconfined. The AdS/CFT correspondence relates $N$ to $k$ through the equality
@@
where $M_{P}=2.4\times 10^{15}$ TeV is the 4D reduced Planck scale, as $2M_{5}^{3}=kM_{P}^{2}$. Requiring the 5D gravity theory to be perturbative imposes the bound $N\gtrsim 5$ Agashe:2007zd ; Chen:2014oha .
The fact that both the AdS-S and warped metric solutions coexist implies the existence of two spacetime background phases. In the regime of small backreaction $u\ll 1$ and tiny temperature, the energy gap between the two phases is dominated by the difference $E_{0}=\left|V_{{\rm eff}}(0)-V_{{\rm eff}}(\rho)\right|$ which can be approximated as Megias:2020vek
@@
Moreover, the free energy associated to these two phases presents a barrier between them, implying the existence of a phase transition between the AdS and RS phases Creminelli:2001th . In the dual language, such a phase transition corresponds to a deconfined-confined FOPT.
@@
2.2 Low-energy setup: $\rho\lesssim$ TeV